# Bibliography management: natbib package

### v1

@@ Abstract
We make use of model-independent statistical methods to assess the consistency of three different supernova compilations: Union3, Pantheon+ and DES 5-year supernovae. We expand the available model space of each, using Crossing Statistics, and test the compatibility of each dataset, against the other two. This is done using (I) a Flat $\Lambda$CDM fitting to, and (II) Iterative Smoothing from, one particular dataset, and determining the level of deformation by required to fit the other two. This allows us to test the mutual consistency of the datasets both within the standard model and in the case of some extended model, motivated by features present in a particular dataset. We find that, in both these cases, the data are only consistent with the point in the parameter space corresponding to zero deformation, at around a $2\sigma$ level, with the DES compilation showing the largest disagreement. However, all three datasets are still found to be consistent to within $1-2\sigma$ for some subset of the extended model space implied by the deformations.
@@ 2 Methodology
In this analysis, we consider data from supernovae observations. For a brief, yet comprehensive history of this observable, see [1]. These data are useful for cosmological applications, as supernovae can provide a distance measure for a particular redshift. Supernovae are called standardisable candles since their light curves may be fitted to a template [^1] Not to be confused with our nomenclature of “template function” in the context of Crossing Statistics, see Section 2.1. in a consistent manner to determine their intrinsic luminosity. Thus, they give us access to the luminosity distance at the redshift measured from the host galaxy. The actual observable is the “distance modulus”:
@@
where $m$ is the effective (after standardisation) apparent magnitude, and $M_{B}$ the B-band absolute magnitude, of the standardised supernova light curve. Although we cannot use supernovae to directly constrain $H_{0}$ (due to its degeneracy with $M_{B}$), we do require its value in order to gain access to the expansion history, and for the purposes of this analysis we choose $h\equiv\frac{H_{0}}{100}=0.6766$, from Planck 2018 results [3]. With this fixed, we may constrain parameters relating to the background expansion: $\Omega_{m}$, $\Omega_{\Lambda}$, etc. depending on the chosen cosmological model.
@@ 2.1 Crossing Statistics
In the limit $n\rightarrow\infty$, the Chebyshev polynomials span the interval $[-1,1]$ and, as such, can provide a basis for reconstructing different functional forms in a model-independent manner. The resulting deformations of the template function will have more degrees of freedom, and can be fit to the same data, to determine if there are any features which the data prefer that are not captured by the initial template function. The best fit crossing hyperparameters $C^{n}_{m}$ ($m\leq n$), found through some sampling method, will reveal how strongly the data support these deviations from the template function. In the case that the fiducial template fit already sufficiently captures the features of the data, then the most likely values of the crossing hyperparameters for the $n^{th}$ order crossing function $(C^{n}_{0},C^{n}_{1},...,C^{n}_{n})$ will be $(1,0,...,0)$, which signifies no effective deformation. Since this analysis involves tests for the consistency of the fiducial templates, we refer to these as the “standard values” of the crossing hyperparameters.
@@
Since $\hat{\mu}^{TF}(z)$ must be related to the actual observed quantity $m$ through the parameter $M_{B}$, which depends on the observation in question, with no direct cosmological significance, there is also a translational degree of freedom in the deformed function. This is not quite the same as the $0^{th}$ order crossing function; while the former provides an additive shift in the estimate for $\mu$, the latter signifies a ‘tilt’ in the function. Thus, at all orders in Crossing Statistics, we allow for $M_{B}$ to vary.
@@ 2.2 Iterative Smoothing
This method, introduced in [24], is broadly a process meant to produce, from data, a function which has a progressively better $\chi^{2}$ fit to that data, in essence completely removing the residuals in the limits of infinite iterations and zero smoothing length. It relies on an initial guess function [^4] In principle, given sufficient iterations, the precise choice of initial function should not matter. that undergoes successive applications by a function that depends on the inverse covariance matrix of the data, the residuals and some weighting function, characterized by a smoothing length, over the domain of the data.
In the case of supernovae distance moduli, the estimate at arbitrary redshift $z$ in the next iteration may be written compactly as follows [26]:
@@ 2.3 Consistency
As discussed previously, we would like to test the mutual consistency of the datasets in two scenarios: that of a model (flat $\Lambda$CDM) and the case where the test is independent of a particular cosmological model (instead there is an implied model which is perhaps an extension beyond the concordance cosmology). In the Bayesian application of Crossing Statistics, we deform a template function by crossing functions in order to provide more freedom and obtain a better fit to the data in a model-independent manner. This additional freedom gives the resulting deformed functions the possibility to be more sensitive to features of the data, which might reveal inconsistencies that go unnoticed when the data are simply fit to a cosmological model.
@@
Both of these null hypotheses are actually twofold. On one hand, we hypothesise that all the datasets are consistent with the individual representation (undeformed template function) of each dataset, under some underlying model (be it $\Lambda$CDM or not); on the other hand, we also hypothesize that there exists a volume within the implied hyperparameter space of a particular representation, where all three datasets are consistent (i.e. could all be produced as realisations from particular deformations of the underlying model). By testing each of these parts of the null hypotheses, we will be able to provide a determination of the mutual consistency of the chosen datasets.
@@ 2.4 Data
  • “Union3” - comprising 2087 SNe Ia collected into 22 bins over the redshift range $2.260.05<z<$ 2.26, standardised and compiled using the Unity 1.5 Bayesian framework [29].
@@
These 3 datasets provide the very necessary empirical observations with which we may do cosmology. The fact that each dataset is a distinct compilation and analysis of supernova data (even though the supernovae of the datasets are not necessarily completely distinct [^5] Over 85% (1363) of the supernovae from Pantheon+ are included in the Union3 compilation, and the 194 low-redshift supernovae from DES-SN5YR are also present in the other two [9].), while supposed to be produced by one underlying model (whether $\Lambda$CDM or otherwise), is ideal leverage to test their consistency with respect to each other.
@@ 3.1 Model-based template function
Table 1: Best fit results for the flat $\Lambda$CDM model parameters from each of the three datasets considered. Note that the much lower value of $\chi^{2}$ for Union3 is due to the fact that the data are binned in 23 redshift bins. In the case of Union3 and DES-SN5YR the data are also given with the $M_{B}$ value already taken care of, and so the small shifts found while fitting (referred to here as $\Delta M_{B}$) are very close to 0.
@@ 3.2 Model-independent template function
We employ iterative smoothing to generate a template function for $\mu(z)$ that is independent of the background cosmological model. In Fig. 2 we show an example of this in the case of Union3 data, by plotting the residuals of the function resulting from iterative smoothing with respect to the cosmological model best fit [^6] The keen-eyed reader might notice the asymmetric distribution of the residuals around the zero line corresponding to the flat $\Lambda$CDM best fit. This is not altogether unexpected, given that the binned Union3 data show strong off-diagonal correlations in the covariance matrix.. As may be seen in the case of Union3, and is indeed true of all cases, the iterative smoothing procedure results in an overall improvement in the $\chi^{2}$ compared to the best fit function, since it serves to create a smooth function that passes as close to as many data points as possible. For a fixed smoothing length $\Delta$, this will not necessarily converge on a function that passes through all the data points, since (even with sufficiently many iterations, $N\rightarrow\infty$, each part of the fitting function is controlled by some range of data points around it. Nevertheless, depending on the distribution of the $\mu$ data in redshift, the $\Delta$ and $N$ must be chosen in such a way as to avoid fitting small scale noise without being completely insensitive to the particular features of the dataset.
@@
The resulting iteratively-smoothed functions will be used as template functions in the second part of the Crossing Statistics analysis. Since each one is constructed as a better fit to its own dataset, they will have taken up some characteristics of their dataset, in a model-independent way. The crossing functions which are then applied after smoothing will allow for large-scale fluctuations of each template function that will slightly deform (and in a sense relax) the characteristics from its own dataset, and determine to what extent these peculiarities can remain, or not, in order to be a good fit to the other datasets. This allows us to test the second null hypothesis.
Figure 3: Evolution of the $\Delta\chi^{2}$ with iteration number, taken with respect to the $\chi^{2}$ of the flat $\Lambda$CDM best fit of each dataset. We see that, although the improvement in $\chi^{2}$ for Pantheon+ is negligible (considering the large number of supernovae, and thus degrees of freedom), the improvement is about an order of magnitude more significant in the case of DES-SN5YR, and even more impressive in the case of Union3.
@@ 4.1 $\Lambda$CDM best fit
Here we show the results of the application of Crossing Statistics, using the flat $\Lambda$CDM best fits to each of the three datasets as the template functions. The likelihoods of the deformed functions of $\mu(z)$, generated by sampling in the space of crossing hyperparameters, and $M_{B}$, are then calculated for each of the three datasets, to determine the possible shapes of $\mu(z)$ suggested by the data. Using the $\chi^{2}$ calculated in this manner, we may determine how the deformations at various orders might improve the fit to the data, and the effect of using the best fit from a different dataset.
In each plot, ‘$C^{n=n}_{i}$’ refers to the case where the template function for angular diameter distance is deformed by the Chebyshev polynomials up to order $n$, when the best fit function is treated as the distance modulus $\mu$. In the case where we only have the template function no crossing functions are used, but the absolute magnitude $M_{B}$ is varied as a nuisance parameter. [^7] Note: Absolute magnitude is not completely degenerate with $C_{0}$. $M_{B}$ is a translation (additive) in $\mu$, while $C_{0}$ is a multiplicative factor. As an example, we show how the application of crossing functions deforms the Pantheon+ best fit template function to better fit the Union3 data.
@@
At each increasing order in Crossing Statistics, the deformed template function is allowed more freedom, and so there is a necessary improvement in the $\chi^{2}$ compared to lower orders (and the template function). For a summary of the initial and best $\chi^{2}$ values, see Table 2. In this table, we also introduce for the first time our abbreviations: for the datasets, to which the deformed functions are fit, we use “P+”, “U3”, “D5” for Pantheon+, Union3 and DES-SN5YR respectively, and we abbreviate ’template function’ to “T.F.”. In Fig. 5, we see how the $\Delta\chi^{2}$ evolves with maximum order of the applied crossing functions. The difference in $\chi^{2}$ of a particular deformed function is shown with respect to the $\chi^{2}$ of the undeformed, template function. It is also important to consider the relative $\Delta\chi^{2}$ that may be achieved by simply using the template function of the data itself (without application of Crossing Statistics), and these are shown as a dashed horizontal lines, coloured according to the corresponding dataset used to generate the template function.
@@
The results from Union3 and Pantheon+ template functions ($\mu^{\rm TF}$) show that, even at the $0^{th}$ order in Crossing Statistics (i.e. a deformation according to $(\mu^{TF}(z)+M_{B})\times C_{0}$), the deformed function can outperform the Flat $\Lambda$CDM best fit from the data, even in the case where the best fit template function does not come from the data itself. This would seem to indicate that the differences between the datasets are not so significant as to prevent the freedoms in one dataset from being explored by a best fit constrained on another dataset. This independence of the new $\chi^{2}$ from the deformed function on the dataset used to determine the best fit points towards a good acceptance of these datasets to the others’ flat $\Lambda$CDM best fit.
@@
This is true for all cases, except where the Union3 template function is used in conjunction with the Pantheon+ data. Here, it is only at $1^{st}$ order in Crossing Statistics that the improvement from using the deformed Union3 template function, surpasses that of the best fit from Pantheon+ (with respect to the undeformed Union3 template function). This means that the Union3 template function is not intrinsically as good a fit to the Pantheon+ data, if given only the 2-parameter freedom of: $M_{B}$ and $C_{0}$, whereas the Pantheon+ template function is, when fit the Union3 data [^8] This may simply be due to the fact that the Pantheon+ dataset has many more points in redshift, and so more freedom is required by the Union3 template function to surpass the Pantheon+ best fit..
@@
By contrast, if we consider the DES-SN5YR best fit template function, we see that it requires the flexibility provided by at least 2nd order in crossing functions before it is able to be a better fit to the other two datasets. It is difficult to draw conclusions from the $\chi^{2}$ alone, but this hints that the DES-SN5YR best fit may be slightly less compatible with the other two datasets, for some reason [^9] As in the case of the Union3 template function, this may be a result of differences in the characteristics of the data, such as the redshift range or sparseness of the data points in redshift, but conclusions cannot be drawn from the $\chi^{2}$ value alone..

### v2

@@ Abstract
We make use of model-independent statistical methods to assess the consistency of three different supernova compilations: Union3, Pantheon+ and DES 5-year supernovae. We expand the available model space of each, using Crossing Statistics, and test the compatibility of each dataset, against the other two. This is done using (I) a Flat $\Lambda$CDM fitting to, and (II) Iterative Smoothing from, one particular dataset, and determining the level of deformation by required to fit the other two. This allows us to test the mutual consistency of the datasets both within the standard model and in the case of some extended model, motivated by features present in a particular dataset. We find that, in both these cases, the data are only consistent with the point in the parameter space corresponding to zero deformation (of I and II), at around a $2\sigma$ level, with the DES compilation showing the largest disagreement. However, all three datasets are still found to be consistent to within $1-2\sigma$ with each other.
@@ 2 Methodology
In this analysis, we consider data from supernovae observations. For a brief, yet comprehensive history of this observable, see [1]. These data are useful for cosmological applications, as supernovae can provide a distance measure for a particular redshift. Supernovae are called standardisable candles since their light curves may be fitted to a template [^1] Not to be confused with our nomenclature of “template function” in the context of Crossing Statistics, see Section 2.1. in a consistent manner to determine their intrinsic luminosity. Thus, they give us access to the luminosity distance at the redshift measured from the host galaxy. The actual observable is the magnitude, which may be related to cosmology by the “distance modulus”:
@@
where $m$ is the effective (after standardisation) apparent magnitude, and $M_{B}$ the B-band absolute magnitude, of the standardised supernova light curve. Although we cannot use supernovae to directly constrain $H_{0}$ (due to its degeneracy with $M_{B}$), we can constrain the combination and, in doing so, gain access to the expansion history. For the purposes of this analysis we report the value of $M_{B}$ obtained when taking $h\equiv\frac{H_{0}}{100km/s/Mpc}=0.6766$, from Planck 2018 results [3]. With this combination fit, we may use Eq. (2.1) to constrain parameters relating to the background expansion: $\Omega_{m}$, $\Omega_{\Lambda}$, etc. depending on the chosen cosmological model.
@@ 2.1 Crossing Statistics
In the limit $n\rightarrow\infty$, the Chebyshev polynomials span the interval $[-1,1]$ and, as such, can provide an orthogonal basis for reconstructing different functional forms in a model-independent manner. The resulting deformations of the template function will have more degrees of freedom, and can be fit to the same data, to determine if there are any features which the data prefer that are not captured by the initial template function. The best fit crossing hyperparameters $C^{n}_{m}$ ($m\leq n$), found through some sampling method, will reveal how strongly the data support these deviations from the template function. In the case that the fiducial template function already sufficiently captures the features of the data, then we would expect that the values of the crossing hyperparameters for the $n^{th}$ order crossing function $(C^{n}_{0},C^{n}_{1},...,C^{n}_{n})$ should be consistent with the point $(1,0,...,0)$, which signifies no effective deformation. Since this analysis involves tests for the consistency of the fiducial templates, we refer to these as the “standard values” of the crossing hyperparameters.
@@
Since $\hat{\mu}^{TF}(z)$ must be related to the actual observed quantity $m$ through the parameter $M_{B}$ which, while having one true underlying value, may have a different inferred value, depending on the subset of data observed, with no direct cosmological significance, there is also a translational degree of freedom in the deformed function. This is not quite the same as the $0^{th}$ order crossing function; while the former provides an additive shift in the estimate for $\mu$, the latter signifies a ‘tilt’ in the function. Thus, at all orders in Crossing Statistics, we allow for $M_{B}$ to vary.
@@ 2.2 Iterative Smoothing
This method, introduced in [24], is broadly a process meant to produce, from data, a function which has a progressively better $\chi^{2}$ fit to that data, in essence completely removing the residuals in the limits of infinite iterations and zero smoothing length. It relies on an initial guess function [^4] In principle, given sufficient iterations, the precise choice of initial function should not matter. that undergoes successive applications by a function that depends on the inverse covariance matrix of the data, the residuals and some weighting function, characterized by a smoothing length, over the domain of the data. In this case, we use the $\Lambda$CDM best fit as the $0^{th}$ iteration of the smoothed function for a particular dataset.
In the case of supernovae distance moduli, the estimate at arbitrary redshift $z$ in the next iteration may be written compactly as follows [28, 29]:
@@ 2.3 Consistency
As discussed previously, we would like to test the mutual consistency of the datasets in two scenarios: that of a model (flat $\Lambda$CDM) and the case where the test is independent of a particular cosmological model (instead there is an implied model which is perhaps an extension beyond the concordance cosmology). In the Bayesian application of Crossing Statistics, we deform a template function by crossing functions in order to provide more freedom and obtain a lower $\chi^{2}$ statistic for the data in a model-independent manner. We choose the $\chi^{2}$ as the statistic of interest here, as an unbiased measure which is simple to calculate, given the standard practice and available likelihood codes for the data, even though the true connection to likelihood is not so straightforward. This additional freedom gives the resulting deformed functions the possibility to be more sensitive to features of the data, which might reveal inconsistencies that go unnoticed when the data are simply fit to a cosmological model.
@@
Both of these null hypotheses are actually twofold. On one hand, we hypothesise that all the datasets are consistent with the individual representation (undeformed template function) of each dataset, under some underlying model (be it $\Lambda$CDM or not); on the other hand, we also hypothesize that there exists a volume within the implied hyperparameter space of a particular representation, where all three datasets are consistent (i.e. could all be produced as realisations from particular deformations of the underlying model). By testing each of these parts of the null hypotheses, we will be able to provide a determination of the mutual consistency of the chosen datasets; we will thus consider the level of consistency by examining the $\sigma$ separations, in the extended parameter space, between the undeformed template function and each dataset’s posterior.The statistical interpretation is as follows: If two dataset contours overlap at, for example, $2-\sigma$, it means that there exists a region in the functional space (corresponding to a possible universe expansion history) which, given the corresponding model, has a $68-95\%$ probability of realising both datasets. The threshold number of $\sigma$ is somewhat arbitray, and we leave it to the reader to impose, based on the required level of stringency.
@@ 2.4 Data
  • “Union3” - comprising a quadratic spline interpolation of 2087 SNe Ia, around a flat $\Lambda$CDM model, at 22 nodes over the redshift range $2.260.05<z<$ 2.26, standardised and compiled using the Unity 1.5 Bayesian framework [31].
@@
The particular data we use are based on the publicly-available likelihood codes, implemented in cobaya and include magnitudes at various redshifts, along with covariance matrices. These 3 datasets provide the very necessary empirical observations with which we may do cosmology. Once again, we draw attention to the fact that Type Ia Supernovae undergo standardisation and fitting, according to a model, so they are only empirical in so far as this process is understood to have been done. The fact that each dataset is a distinct compilation and analysis of supernova data (even though the supernovae of the datasets are not necessarily completely distinct [^5] Over 85% (1363) of the supernovae from Pantheon+ are included in the Union3 compilation, and the 194 low-redshift supernovae from DES-SN5YR are also present in the other two [9].), while supposed to be produced by one underlying model (whether $\Lambda$CDM or otherwise), is ideal leverage to test their consistency with respect to each other.
@@ 3.1 Model-based template function
$\chi^{2}/DoF$ 0.912 1.1980 0.946
Table 1: Best fit results for the flat $\Lambda$CDM model parameters from each of the three datasets considered. Note that the much lower value of $\chi^{2}$ for Union3 is due to the fact that the data are interpolated $2^{\rm nd}$ order splines around a flat $\Lambda$CDM model, at 22 node points. In the case of Union3 and DES-SN5YR the data are also given with the $M_{B}$ value already taken care of, and so the small shifts found while fitting (referred to here as $\Delta M_{B}$) are very close to 0. Here “DoF” refers to the number of degrees of freedom, which is the (effective) number of data points in each dataset minus the number of parameters fit for a particular model [30, 31, 32].
@@ 3.2 Model-independent template function
We employ iterative smoothing to generate a template function for $\mu(z)$ that is independent of the background cosmological model, taking as our $0^{th}$ iteration, the $\Lambda$CDM best fit from each dataset. In Fig. 2 we show an example of this in the case of Union3 data, by plotting the residuals of the function resulting from iterative smoothing with respect to the cosmological model best fit [^6] The keen-eyed reader might notice the asymmetric distribution of the residuals around the zero line corresponding to the flat $\Lambda$CDM best fit. This is not altogether unexpected, given that the splined Union3 data show strong off-diagonal correlations in the covariance matrix.. As may be seen in the case of Union3, and is indeed true of all cases, the iterative smoothing procedure results in an overall improvement in the $\chi^{2}$ compared to the best fit function (see Fig. 3), since it serves to create a smooth function that passes as close to as many data points as possible. For a fixed smoothing length $\Delta$, this will not necessarily converge on a function that passes through all the data points, since there are only two parameters ($\Delta$ and $N$). Even with sufficiently many iterations, $N\rightarrow\infty$, each part of the fitting function is controlled by some range of data points around it. Nevertheless, depending on the distribution of the $\mu$ data in redshift, the $\Delta$ and $N$ must be chosen in such a way as to avoid fitting small scale noise without being completely insensitive to the particular features of the dataset.
@@
The resulting iteratively-smoothed functions will be used as template functions in the second part of the Crossing Statistics analysis. Since each one is constructed to more closely fit to its own dataset, they will have taken up some characteristics of their dataset, in a model-independent way. The crossing functions which are then applied after smoothing will allow for large-scale fluctuations of each template function that will slightly deform (and in a sense relax) the characteristics from its own dataset, and determine to what extent these peculiarities can remain, or not, in order to be a good fit to the other datasets. This allows us to test the second null hypothesis.
Figure 3: Evolution of the $\Delta\chi^{2}$ with iteration number, taken with respect to the $\chi^{2}$ of the flat $\Lambda$CDM best fit of each dataset. We see that, although the improvement in $\chi^{2}$ for Pantheon+ is negligible (considering the large number of supernovae, and thus degrees of freedom), the improvement is about an order of magnitude more significant in the case of DES-SN5YR. The iterative smoothing function successfully converges on a closer fit to the data, and thus represents a reasonable model-independent function through the data.
@@ 4.1 $\Lambda$CDM best fit
Here we show the results of the application of Crossing Statistics, using the flat $\Lambda$CDM best fits to each of the three datasets as the template functions. The posteriors of the deformed functions of $\mu(z)$ (generated by MCMC sampling in the space of crossing hyperparameters, and $M_{B}$) are then calculated for each of the three datasets, to determine the possible shapes of $\mu(z)$ suggested by the data. Using the $\chi^{2}$ calculated in this manner, we may determine how the deformations at various orders might improve the fit to the data, and the effect of using the best fit from a different dataset.
In Fig. 4, ‘$C^{n=k}_{i}$’ refers to the case where the template function for distance modulus is deformed by the Chebyshev polynomials up to order $k$. In the case where we only have the template function no crossing functions are used, but the absolute magnitude $M_{B}$ is varied as a nuisance parameter. [^7] Note: Absolute magnitude is not completely degenerate with $C_{0}$. $M_{B}$ is a translation (additive) in $\mu$, while $C_{0}$ is a multiplicative factor. As an example, we show how the application of crossing functions deforms the Pantheon+ best fit template function to better fit the Union3 data.
@@
At each increasing order in Crossing Statistics, the deformed template function is allowed more freedom, and so there is a necessary improvement in the $\chi^{2}$ compared to lower orders (and the template function). For a summary of the initial and best $\chi^{2}$ values, see Table 2. In this table, we also introduce for the first time our abbreviations: for the datasets, to which the deformed functions are fit, we use “P+”, “U3”, “D5” for Pantheon+, Union3 and DES-SN5YR respectively, and we abbreviate ’template function’ to “T.F.”.
In Fig. 5, we see how the $\Delta\chi^{2}$ evolves with maximum order of the applied crossing functions, from “No Crossing” (undeformed template function) to $4^{\rm th}$ order (with 5 additional free parameters). This is calculated as $\Delta\chi^{2}=\chi^{2}\left[\hat{\mu}^{TF,A}(z)\sum_{i=0}^{n}C^{n}_{i}T_{i}(\tilde{z})\right]_{B}-\chi^{2}\left[\hat{\mu}^{TF,A}(z)\right]_{B}$, i.e. the difference in $\chi^{2}$ between a deformed template function A, fit to dataset B, and the undeformed version of the same template function A, fit to the same dataset. In this sense, it shows the decrease in $\chi^{2}$ as more freedom [^8] For a discussion concerning the degrees of freedom and over-fitting, see Appendix A is given to the deformation, by increasing the number of hyperparameters. For each panel, the undeformed template function A is fixed, and stated on the y-axis. The curves in a particular panel then correspond to the deformations (at various orders, stated on the x-axis) of this template function, to each each dataset. In the “No Crossing” case corresponds to the undeformed template function, it is always identically zero and is shown for reference.
It is also important to consider the relative $\Delta\chi^{2}$ that may be achieved by simply using the template function of the data itself (without application of Crossing Statistics), i.e. $\Delta\chi^{2}=\chi^{2}\left[\hat{\mu}^{TF,B}(z)\right]_{B}-\chi^{2}\left[\hat{\mu}^{TF,A}(z)\right]_{B}$. These are shown as a dashed horizontal lines, coloured according to the corresponding dataset used to generate the template function.
@@
The results (in the left and middle panels) from Union3 and Pantheon+ template functions ($\mu^{\rm TF}$) show that, even at the $0^{th}$ order in Crossing Statistics (i.e. a deformation according to $(\mu^{TF}(z)+M_{B})\times C_{0}$), the deformed function can outperform the Flat $\Lambda$CDM best fit from the data, even in the case where the best fit template function does not come from the data itself. This would seem to indicate that the differences between the datasets are not so significant as to prevent the freedoms in one dataset from being explored by a best fit constrained on another dataset. This independence of the new $\chi^{2}$ from the deformed function on the dataset used to determine the best fit points towards a good acceptance of these datasets to the others’ flat $\Lambda$CDM best fit.
@@
This is true for all these cases, except where the Union3 template function is used in conjunction with the Pantheon+ data (middle panel, blue curve). Here, it is only at $1^{st}$ order in Crossing Statistics that the improvement from using the deformed Union3 template function, surpasses that of the best fit from Pantheon+ (with respect to the undeformed Union3 template function). This means that the Union3 template function is not intrinsically as good a fit to the Pantheon+ data, if given only the 2-parameter freedom of: $M_{B}$ and $C_{0}$, whereas the Pantheon+ template function is, when fit the Union3 data [^9] This may simply be due to the fact that the Pantheon+ dataset has many more points in redshift, and so more freedom is required by the Union3 template function to surpass the Pantheon+ best fit..
@@
By contrast, if we consider the DES-SN5YR best fit template function (right panel), we see that it requires the flexibility provided by at least 2nd order in crossing functions before it is able to obtain a lower $\chi^{2}$ in the context of the other two datasets (pink and blue curves). It is difficult to draw conclusions from the $\chi^{2}$ alone, but this hints that the DES-SN5YR best fit may be slightly less compatible with the other two datasets, for some reason [^10] As in the case of the Union3 template function, this may be a result of differences in the characteristics of the data, such as the redshift range or sparseness of the data points in redshift, but conclusions cannot be drawn from the $\chi^{2}$ value alone..